# Secondary spins of extreme mass ratio inspirals: Measurement precision and astrophysical applications

### v1

@@ Abstract
Extreme mass-ratio inspirals (EMRIs), consisting of a secondary (stellar mass) black hole (BH) orbiting around a supermassive BH, are one of the primary targets for future spaceborne gravitational wave (GW) detectors. The spin of the secondary BH encodes the formation history of the stellar mass BH and the formation process of the EMRI. In this work, we construct a kludge EMRI waveform model taking the secondary spin into account and forecast the measurement precision of the secondary spin by future spaceborne GW detectors with the Fisher information matrix. We find the secondary spin should be measured with a reasonably good precision for generic eccentric and inclined EMRIs, though the measurement precision largely degrades for low-eccentricity and nearly equatorial EMRIs. As an example of its astrophysical applications, we propose that the secondary spin can be used for distinguishing dry (loss cone) EMRIs (where the secondary BHs were born in the collapse of individual massive stars and are of low spin) and Hills EMRIs (where the secondary BHs are remnants of massive star binaries and the secondary spins follow a bimodal distribution).
@@ I Introduction
The EMRI GW signals in the sensitivity band of spaceborne GW detectors in general last for a few months to a few years depending on the masses of the system, and the compact object completes $\sim 10^{4}-10^{5}$ cycles around the SMBH. As a result, the intrinsic EMRI parameters (e.g., the mass and spin of SMBH, the mass of the secondary object) can be measured to a superb precision, with relative uncertainty as low as $\sim 10^{-4}-10^{-6}$ Babak et al. (2017a), making EMRI a unique tool to resolve the nature of the SMBH and test general relativity Gair et al. (2013). EMRIs are also well known as a sensitive probe to the SMBH environment, including a nearby third body Bonga et al. (2019), the accretion disk in the case of accreting SMBH Kocsis et al. (2011); Yunes et al. (2011); Speri et al. (2023); Chatterjee et al. (2023); Duque et al. (2024a) and a possible axion-like cloud Zhang & Yang (2020, 2019); Brito & Shah (2023); Duque et al. (2024b) . In this paper, we mainly focus on the spin of the secondary body, the measurement precision and the astrophysical applications, which are largely unexplored.
@@
Generally, the dynamic of a spinning particle in a curved spacetime is described by so-called Mathison-Papapetrous-Dixon (MPD) equations, which are proposed and developed by Mathisson Mathisson (1937) and Papapetrou Papapetrou (1951) in order to solve the extended body problem in general relativity, and then reformulated by Dixon Dixon (1970, 1974). As for the EMRI system, the effect of the secondary spin emerges at the first order of mass ratio, in the form of spin-curvature force following the geodesic equation. Some special orbits have been carefully studied, such as the straight falling one in Ref. Mino et al. (1996), and the circular-aligned one in Ref. Tanaka et al. (1996); Han (2010); Huerta & Gair (2011); Huerta et al. (2012); Piovano et al. (2021); Jiang & Han (2024); Jiang et al. (2024); Rahman & Bhattacharyya (2023). So far, there is no reliable waveform model to describe the generic case, but a lot of efforts have been made recently. Ref. Drummond & Hughes (2022) computes the bound orbit by associating it with a reference geodesic which shares the same turning points. Ref. Drummond et al. (2024) develops a model that treats the generic inspiral as a sequence of geodesic orbits. And recently, Ref. Skoupý & Witzany (2024) introduces a coordinate shift of a generic orbit to the virtual geodesic worldline in the linear-spin approximation, which leads to the separable formula and thus should be useful for fully relativistic waveform modeling.
@@
In Ref. Huerta & Gair (2011), the authors develop a kludge waveform model, which uses the MPD-based equations of motion derived by Ref. Saijo et al. (1998) in an equatorial-aligned case for spinning black hole binaries, and includes first-order conservative self-force corrections to compute the evolution of the inspiraling sBH’s orbital frequency. Their analysis demonstrates that LISA is unable to measure the spin of sBH inspiraling into the Kerr SMBH with a typically mass-ratio $\sim 10^{-5}$, as the mean error is much larger than 1. Some follow-up studies confirm this conclusion, such as Ref. Huerta et al. (2012) which improves the waveform model by extending to higher-order spin effects, Refs. Piovano et al. (2021) which uses fully relativistic numerical waveforms to the leading order in an adiabatic expansion, and Ref. Burke et al. (2024) which employs Bayesian inference to assess the importance of first post-adiabatic terms. However, in these researches mentioned above, the motion of the sBH is strictly constrained to an equatorial and further circular orbit, where the secondary spin must be aligned to the primary one. This constrained case may not represent the full picture of the secondary spin in real EMRI dynamics. As we will show later, the secondary spin plays a much more important role in EMRI dynamics for generic inclined and eccentric orbits, therefore will be more tightly constrained from EMRI waveforms.
@@
Actually, the secondary spin in the EMRI system contains rich astrophysical information about the sBH formation and the EMRI formation processes. As is studied with LVK binary BH mergers in Ref. Belczynski et al. (2020), the spin of sBH combined with the mass can reveal the evolutionary road (such as angular momentum transport) of massive stars. Similar to LVK binary BH mergers, a number of EMRI formation channels have been proposed. The most considered EMRI sources for LISA are the dry (loss-cone) EMRIs Hopman & Alexander (2005); Preto & Amaro-Seoane (2010); Bar-Or & Alexander (2016); Babak et al. (2017b); Amaro-Seoane (2018); Broggi et al. (2022), in which a stellar compact object is captured by SMBH by being pushed into a gravitational radiation-efficient orbit via multi-body scatterings within the galactic nuclei. In addition, Hills EMRIs proposed by Ref. Hills (1988); Miller et al. (2005), in which one of the binary BHs is captured due to tidal disruption near the SMBH, should also contribute a fraction of detectable EMRIs. Another channel that contributes comes from the SMBH associated with an accretion disk, as known in active galactic nuclei (AGN), the motion of sBH is dampened in the interaction with the disk, finally captured onto the disk plane and migrates inward in the accretion disk, and thus is named wet EMRIs Sigl et al. (2007); Levin (2007); Pan & Yang (2021a); Pan et al. (2021); Pan & Yang (2021b); Pan et al. (2022); Derdzinski & Mayer (2023).
@@
A key question that remains to be answered is: what is the unique signature of each formation channel that is observable in future EMRI detections. A clear signature of wet EMRIs is the low orbital eccentricity, which is distinct from EMRIs in the other two channels. Hills EMRIs were previously thought to be circular in the LISA sensitivity band Miller et al. (2005), therefore are obviously different from dry EMRIs. But more recent detailed simulations show that dry and Hills EMRIs in the mHz band are actually quite similar in their orbital properties Raveh & Perets (2021): high orbital eccentricities and no preferred orbital orientation. As a piece of indirect evidence for different EMRI formation channels, recent analyses of quasi-periodic eruptions, which are likely EM counterparts to low-frequency EMRIs with orbital frequency $f_{\rm obt}\sim 10^{-5}$ Hz, show that there are indeed clues to two different populations of EMRIs Zhou et al. (2024, 2024b, 2024a): a low-eccentricity population that is consistent with the wet channel prediction and a high-eccentricity population that is consistent with both the dry and the Hills channels. It is still unclear how to distinguish dry EMRIs from Hills EMRIs due to their similarity in orbital properties. In this work, we point out that the secondary spin can be used as a discriminator: the dry EMRIs are of low secondary spins due to the efficient angular momentum transport from the core to the envelope in the late evolution stage of massive stars as informed by LVK BH spins Abbott et al. (2019); Roulet & Zaldarriaga (2019); Fuller & Ma (2019); Abbott et al. (2023), while the Hills EMRIs are expected to be double peaked in secondary spins, because the second born BH in a BH binary is expected to spin faster due to the tidal spin up in the Helium core stage by the first born BH Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel & Fragos (2020).
@@
In this paper, we construct a kludge EMRI waveform model taking the secondary spin into account by numerically integrating MPD equations while ignoring the effect of the secondary spin in radiation reactions. With this waveform model, we then forecast the measurement precision of secondary spins by Taiji. We find the secondary spin should be measured to a reasonable precision for generic inclined and eccentric orbits while almost unconstrained for equatorial and circular orbits. As an astrophysical application, we propose that the secondary spins can be used for distinguishing dry and Hills EMRIs: dry EMRIs are expected to be of low secondary spins, and a robust detection of high secondary spins will be a clear signature of Hills EMRIs. Note that Taiji and LISA are pretty the same in the designed sensitivity (see Fig. 1), so the forecast result in this work also equally applies to LISA.
@@
This paper is organized as follows. In Sec. II.1, we start with introducing the waveform model used in this paper, and then in Sec. II.2, we introduce the data analysis method and we present the forecast measurement precision of the secondary spin. In Sec. III.1, we introduce the main feature of the three EMRI channels, then propose a distribution of the secondary spin for dry and Hills EMRIs. Finally, in Sec. III.2 and Sec. III.3, we present the detailed inference result of formation channels (the recovered distribution). We conclude this work with Sec. IV. Throughout this paper, the geometrical units with $c=G=1$ and the Einstein summation convention are adopted.
@@ II.1 Waveform model
$\frac{dp^{\mu}}{d\tau}$ $=-\frac{1}{2}S^{\kappa\lambda}v^{\rho}R^{\mu}_{\ \rho\kappa\lambda}-\Gamma^{\mu}_{\kappa\lambda}p^{\kappa}v^{\lambda},$ (1)
$\frac{dS^{\mu\nu}}{d\tau}$ $=2p^{[\mu}v^{\nu]}-\Gamma^{\mu}_{\kappa\lambda}S^{\kappa\nu}v^{\lambda}-\Gamma^{\nu}_{\kappa\lambda}S^{\mu\kappa}v^{\lambda},$ (2)
$mv^{\mu}$ $=p^{\mu}+\frac{2S^{\mu\nu}R_{\nu\kappa\lambda\rho}p^{\kappa}S^{\lambda\rho}}{4+R_{\kappa\lambda\rho\sigma}S^{\kappa\lambda}S^{\rho\sigma}},$ (3)
@@
along with $p^{\mu}v_{\mu}=-m$. Apart from $S^{\mu\nu}$, another frequently used spin quantity is the spin vector $S^{\mu}$ which can be expressed as
@@
Meanwhile, we can get a constant of motion if the spacetime admits symmetries described by a Killing vector field $\xi^{\alpha}$ Rüdiger & Penrose (1981),
@@
And in Kerr spacetime, a quadratic “conserved” quantity $Q$, which corresponds to the known Carter constant, can be deduced through linear spin approximation Rudiger (1983); Compère & Druart (2022), which takes the form
@@
where the post-Newtonian fluxes $\left\{\dot{E},\dot{L_{z}},\dot{Q}\right\}$ in terms of the orbit elements $\left\{p,e,\iota\right\}$ we adopted are those in Ref. Gair & Glampedakis (2006). During inspirals, semi-latus rectum $p$, eccentricity $e$ and “inclination angle” $\iota$ are calculated through
@@
For evolving the equations of motion above [Eqs. (23,15)], we need to specify the initial values of dynamical variables $\left\{S^{\mu},p^{\mu},x^{\mu}\right\}$. In the case of generic misaligned and eccentric orbits, we choose
@@
as the initial intrinsic parameters, while initial time $t_{0}$ and initial phase in the azimuthal direction $\phi_{0}$ are set to 0, and $s$ is the dimensionless spin parameter of the sBH which is defined as $s\equiv\frac{1}{m}\sqrt{S^{\mu}S_{\mu}}$. Then $\left\{p^{\mu}_{0}\right\}$ are resolved by the relationship between orbit elements and momenta for geodesic orbits Schmidt (2002), and $\left\{S^{t}_{0},S^{\theta}_{0}\right\}$ are resolved by Eq. (6).
@@
LISA and Taji will detect the GW signal with an interferometric measurement of differential optical pathlength modulation along the arms between three free-falling test masses inside spacecraft, i.e. the relative frequency shift $\left\{y_{12}(t),y_{21}(t),y_{13}(t),y_{31}(t),y_{23}(t),y_{32}(t)\right\}$ where the lower indices mark the spacecraft. A change of light-travel distance caused by the GW should be computed by the projection of the above SSB-frame waveforms onto the detector arms, and the response function depends on not only the frequency of signals but also the time due to the typically long duration of EMRI waveforms. Moreover, the laser noise will drown out the GW signal with realistically unequal arms, thus time-delay interferometry (TDI) technology is proposed by Ref. Tinto & Armstrong (1999); Tinto et al. (2002, 2004); Tinto & Dhurandhar (2014). In short, the key point of TDI is the linear combination of time-delay GW projections in order to construct equivalent equal-length arms and suppress the laser noise. For the second-generation TDI designed for evolving length of arms, the Michelson combination, $X$, is given by Tinto et al. (2004)
@@
Other intrinsic parameters are set as $m=10M_{\odot}$, $M=1\times 10^{6}M_{\odot}$, $a=0.99M$, $S^{r}_{0}=S^{\phi}_{0}=0$, $p_{0}=10M$, $e_{0}=0.2$, $\iota_{0}=0.5$, $r_{0}=p_{0}$, $\theta_{0}=\frac{\pi}{2}$ (where the subscribe “0” means the initial time), and extrinsic parameters are set as $T=1\text{yr}$, $\theta_{A}=\frac{\pi}{2}$, $\phi_{A}=\frac{\pi}{5}$, $\theta_{N}=\frac{\pi}{3}$, $\phi_{N}=\frac{\pi}{4}$, $\Phi=\frac{\pi}{2}$, $D=1$ Gpc. Note that, although we take initial radial and azimuthal components of $S^{\mu}$ to be 0, they will become nonzero as a result of precession for generic misaligned orbits.
Figure 1: The characteristic strain $h_{c}$ for $s=0.1$ (blue curve) and $s=0.9$ (purple curve), and the noise amplitude $h_{n}$ of Taiji (lighgrey curve) and LISA (black curve).

### v3

# Secondary spins of extreme mass-ratio inspirals: A probe to the formation channels

@@ Abstract
Extreme mass-ratio inspirals (EMRIs), consisting of a secondary (stellar mass) black hole (BH) orbiting around a supermassive BH, are one of the primary targets for future spaceborne gravitational wave (GW) detectors. The spin of the secondary BH encodes the formation history of the stellar mass BH and the formation process of the EMRI. In this work, we construct a kludge EMRI waveform model taking the secondary spin into account and preliminarily forecast the measurement precision of the secondary spin by future spaceborne GW detectors with the Fisher information matrix. We find the secondary spin might be measured with reasonably good precision for generic eccentric and inclined EMRIs, with the caveat that the predictive precision may be constrained by the model’s inherent simplifications. As an example of its astrophysical applications, we propose that the secondary spin can be used for distinguishing dry (loss cone) EMRIs (where the secondary BHs were born in the collapse of individual massive stars and are of low spin) and Hills EMRIs (where the secondary BHs are remnants of massive star binaries and the secondary spins follow a bimodal distribution).
@@ I Introduction
The EMRI GW signals in the sensitivity band of spaceborne GW detectors in general last for a few months to a few years depending on the masses of the system, and the compact object completes $\sim 10^{4}-10^{5}$ cycles around the SMBH. As a result, the intrinsic EMRI parameters (e.g., the mass and spin of SMBH, the mass of the secondary object) can be measured to a superb precision, with relative uncertainty as low as $\sim 10^{-4}-10^{-6}$ Babak et al. (2017a), making EMRI a unique tool to resolve the nature of the SMBH and test general relativity Gair et al. (2013). EMRIs are also well known as a sensitive probe to the SMBH environment Dyson et al. (2025), including a nearby third body Bonga et al. (2019), the accretion disk in the case of accreting SMBH Kocsis et al. (2011); Yunes et al. (2011); Speri et al. (2023); Chatterjee et al. (2023); Duque et al. (2024a) and a possible axion-like cloud Zhang and Yang (2020, 2019); Brito and Shah (2023); Duque et al. (2024b) . In this paper, we mainly focus on the spin of the secondary body, the measurement precision and the astrophysical applications, which are largely unexplored.
@@
Generally, the dynamic of a spinning particle in a curved spacetime is described by so-called Mathison-Papapetrous-Dixon (MPD) equations, which were proposed and developed by Mathisson Mathisson (2010) and Papapetrou Papapetrou (1951) in order to solve the extended body problem in general relativity, and then reformulated by Dixon Dixon (1970, 1974). As for the EMRI system, the effect of the secondary spin emerges at the first order of mass ratio, in the form of spin-curvature force following the geodesic equation. Some special orbits have been carefully studied, such as the straight falling one in Ref. Mino et al. (1996), and the circular-aligned one in Ref. Tanaka et al. (1996); Han (2010); Huerta and Gair (2011); Huerta et al. (2012); Piovano et al. (2021); Jiang and Han (2024); Jiang et al. (2024); Rahman and Bhattacharyya (2023). So far, there is no well-established waveform template to describe the generic case, but a lot of efforts have been made recently. Ref. Drummond and Hughes (2022) computed the bound orbit by associating it with a reference geodesic that shares the same turning points. Ref. Skoupý et al. (2023) calculated the asymptotic GW energy and angular momentum fluxes for generic orbits in the linear spin approximation, and Ref. Piovano et al. (2024) improved it by using the Hamilton-Jacobi formalism found in Ref. Witzany (2019). Ref. Skoupý and Witzany (2024a) calculated the post-Newtonian (PN) expansion of eccentric orbits of spinning bodies around Schwarzschild black holes and then computed the PN fluxes from these orbits. Ref. Drummond et al. (2024) developed a model that treats the generic inspiral as a sequence of geodesic orbits. And recently, Ref. Skoupý and Witzany (2024b) introduced a coordinate shift of a generic orbit to the virtual geodesic worldline in the linear-spin approximation, which leads to the separable formula and thus should be useful for fully relativistic waveform modeling.
@@
In Ref. Huerta and Gair (2011), the authors developed a kludge waveform model, which uses the MPD-based equations of motion derived by Ref. Saijo et al. (1998) in an equatorial-aligned case for spinning black hole binaries and includes first-order conservative self-force corrections to compute the evolution of the inspiraling sBH’s orbital frequency. Their analysis demonstrates that LISA is unable to measure the spin of sBH inspiraling into the Kerr SMBH with a typically mass-ratio $\sim 10^{-5}$, as the mean error is much larger than 1. Some follow-up studies confirm this conclusion, such as Ref. Huerta et al. (2012) which improves the waveform model by extending to higher-order spin effects, Refs. Piovano et al. (2021) which uses fully relativistic numerical waveforms to the leading order in an adiabatic expansion, and Ref. Burke et al. (2024) which employs Bayesian inference to assess the importance of first post-adiabatic terms. However, in these researches mentioned above, the motion of the sBH is strictly constrained to an equatorial and further circular orbit, where the secondary spin must be aligned to the primary one. This constrained case may not represent the full picture of the secondary spin in real EMRI dynamics. As we will show later, the secondary spin plays a much more important role in EMRI dynamics for generic inclined and eccentric orbits, therefore will be more tightly constrained from EMRI waveforms.
@@
Actually, the secondary spin in the EMRI system contains rich astrophysical information about the sBH formation and the EMRI formation processes. As was studied with LVK binary BH mergers in Ref. Belczynski et al. (2020), the spin of sBH combined with the mass can reveal the evolutionary road (such as angular momentum transport) of massive stars. Beside, Ref. Piovano et al. (2020); Xu et al. (2025) suggested that measurements of the secondary spin may be used to distinguish the sBH from other compact objects. Similar to LVK binary BH mergers, a number of EMRI formation channels have been proposed. The most considered EMRI sources for LISA are the dry (loss-cone) EMRIs Hopman and Alexander (2005); Preto and Amaro-Seoane (2010); Bar-Or and Alexander (2016); Babak et al. (2017b); Amaro-Seoane (2018); Broggi et al. (2022), in which a stellar compact object is captured by SMBH by being pushed into a gravitational radiation-efficient orbit via multi-body scatterings within the galactic nuclei. In addition, Hills EMRIs proposed by Ref. Hills (1988); Miller et al. (2005), in which one of the binary BHs is captured due to tidal disruption near the SMBH, should also contribute a fraction of detectable EMRIs. Another channel that contributes comes from the SMBH associated with an accretion disk, as known in active galactic nuclei (AGN), the motion of sBH is dampened in the interaction with the disk, finally captured onto the disk plane and migrates inward in the accretion disk, and thus is named wet EMRIs Sigl et al. (2007); Levin (2007); Pan and Yang (2021a); Pan et al. (2021); Pan and Yang (2021b); Pan et al. (2022); Derdzinski and Mayer (2023).
@@
A key question that remains to be answered is: What is the unique signature of each formation channel that is observable in future EMRI detections? A clear signature of wet EMRIs is the low orbital eccentricity, which is distinct from EMRIs in the other two channels. Hills EMRIs were previously thought to be circular in the LISA sensitivity band Miller et al. (2005), therefore are obviously different from dry EMRIs. But more recent detailed simulations show that dry and Hills EMRIs in the mHz band are actually quite similar in their orbital properties Raveh and Perets (2021): high orbital eccentricities and no preferred orbital orientation. As a piece of indirect evidence for different EMRI formation channels, recent analyses of quasi-periodic eruptions, which are likely EM counterparts to low-frequency EMRIs with orbital frequency $f_{\rm obt}\sim 10^{-5}$ Hz, show that there are indeed clues to two different populations of EMRIs Zhou et al. (2024, 2024a, 2024b): a low-eccentricity population that is consistent with the wet channel prediction and a high-eccentricity population that is consistent with both the dry and the Hills channels. It is still unclear how to distinguish dry EMRIs from Hills EMRIs due to their similarity in orbital properties. In this work, we point out that the secondary spin can be used as a discriminator: the dry EMRIs are of low secondary spins due to the efficient angular momentum transport from the core to the envelope in the late evolution stage of massive stars as informed by LVK BH spins Abbott et al. (2019); Roulet and Zaldarriaga (2019); Fuller and Ma (2019); Abbott et al. (2023), while the Hills EMRIs are expected to be double peaked in secondary spins, because the second born BH in a BH binary is expected to spin faster due to the tidal spin up in the Helium core stage by the first born BH Zaldarriaga et al. (2018); Bavera et al. (2020); Mandel and Fragos (2020); Adamcewicz et al. (2024).
@@
In this paper, we construct a kludge EMRI waveform model taking the secondary spin into account by numerically integrating MPD equations and PN radiation reactions. With this waveform model, we then forecast the measurement precision of secondary spins by Taiji. We find the secondary spin should be measured to a reasonable precision for generic inclined and eccentric orbits while almost unconstrained for equatorial and circular orbits. As an astrophysical application, we propose that the secondary spins can be used for distinguishing dry and Hills EMRIs: dry EMRIs are expected to be of low secondary spins, and a robust detection of high secondary spins will be a clear signature of Hills EMRIs. Note that Taiji and LISA are pretty the same in the designed sensitivity (see Fig. 1), so the forecast result in this work also equally applies to LISA.
But, the readers should pay attention to the fact that the EMRI waveform model in this paper is not self-consistent, as we exclude some same-order terms. Specifically, due to the extremely small mass ratio $\epsilon$ in EMRI systems, there exists a known two-timescale expansion LISA Consortium Waveform Working Group et al. (2023). That is to say, the orbit phases $\psi_{A}$ evolve on the orbit time scale $\sim 2\pi/\Omega_{A}$ as
$\frac{d\psi_{A}}{dt}=\Omega^{(0)}_{A}\left(J^{B}\right)+\epsilon\Omega^{(1)}_{A}\left(J^{B}\right)+\mathcal{O}\left(\epsilon^{2}\right)\ ,$ (1)
where $\Omega_{A}$ are the orbit frequencies and $J^{A}$ are the system parameters (such as semi-latus rectum and eccentricity). While $J^{A}$ evolve slowly on the inspiral time scale $\sim 2\pi/(\epsilon\Omega_{B})$ as
$\frac{dJ^{A}}{dt}=\epsilon G^{A}_{(1)}\left(J^{B}\right)+\epsilon^{2}G^{A}_{(2)}\left(J^{B}\right)+\mathcal{O}\left(\epsilon^{3}\right)\ ,$ (2)
where $G^{A}$ are the time-averaged dissipative effects. For compact secondary objects, the spin (finite-size effect) contributes the first-order conservative and the second-order dissipative forces in $\Omega^{(1)}_{A}$ and $G^{A}_{(2)}$, which are often named the Post-1-adiabatic (1PA) terms. The 1PA terms should also include the finite-mass self-force effects, but only the conservative part of spin effects is considered in this paper. This simplified dynamics modeling may overestimate or underestimate the accuracy of the secondary spin measurements, and yet the secondary spins should still be a useful probe for formation channels in future observations.
@@
This paper is organized as follows: In Sec. II.1, we start with introducing the waveform model used in this paper, and then in Sec. II.2, we introduce the data analysis method and we present the forecast measurement precision of the secondary spin. In Sec. III.1, we introduce the main feature of the three EMRI channels, then propose a distribution of the secondary spin for dry and Hills EMRIs. Finally, in Sec. III.2 and Sec. III.3, we present the detailed inference result of formation channels (the recovered distribution). We conclude this work with Sec. IV. Throughout this paper, the geometrical units with $c=G=1$ and the Einstein summation convention are adopted.
@@ II.1 Waveform model
$\frac{dp^{\mu}}{d\tau}$ $=-\frac{1}{2}S^{\kappa\lambda}v^{\rho}R^{\mu}_{\ \rho\kappa\lambda}+\Gamma^{\mu}_{\kappa\lambda}p^{\kappa}v^{\lambda},$ (3)
$\frac{dS^{\mu\nu}}{d\tau}$ $=2p^{[\mu}v^{\nu]}+\Gamma^{\mu}_{\kappa\lambda}S^{\kappa\nu}v^{\lambda}+\Gamma^{\nu}_{\kappa\lambda}S^{\mu\kappa}v^{\lambda},$ (4)
$v^{\mu}$ $=\frac{\mathcal{M}}{m^{2}}\left(p^{\mu}+\frac{2S^{\mu\nu}R_{\nu\kappa\lambda\rho}p^{\kappa}S^{\lambda\rho}}{4m^{2}+R_{\kappa\lambda\rho\sigma}S^{\kappa\lambda}S^{\rho\sigma}}\right),$ (5)
@@
along with $p^{\mu}v_{\mu}=-\mathcal{M}$. Apart from $S^{\mu\nu}$, another frequently used spin quantity is the spin vector $S^{\mu}$ which can be expressed as
@@
Meanwhile, we can get a constant of motion if the spacetime admits symmetries described by a Killing vector field $\xi^{\alpha}$ Rüdiger and Penrose (1981),
@@
And in Kerr spacetime, a quadratic “conserved” quantity $Q$, which corresponds to the known Carter constant, can be deduced through linear spin approximation Rudiger (1983); Compère and Druart (2022), which takes the form
@@
where the post-Newtonian fluxes $\left\{\dot{E},\dot{L_{z}},\dot{Q}\right\}$ in terms of the orbit elements $\left\{p,e,\iota\right\}$ we adopted are those in Ref. Gair and Glampedakis (2006). During inspirals, semi-latus rectum $p$, eccentricity $e$ and “inclination angle” $\iota$ are calculated through
@@
For evolving the equations of motion above [Eqs. (45,17)], we need to specify the initial values of dynamical variables $\left\{S^{\mu},p^{\mu},x^{\mu}\right\}$. In the case of generic misaligned and eccentric orbits, we choose
@@
as the initial intrinsic parameters, while initial time $t_{0}$ and initial phase in the azimuthal direction $\phi_{0}$ are set to 0, and $s$ is the dimensionless spin parameter of the sBH which is defined as $s\equiv\frac{1}{m}\sqrt{S^{\mu}S_{\mu}}$. Then $\left\{p^{\mu}_{0}\right\}$ are resolved by the relationship between orbit elements and momenta for geodesic orbits Schmidt (2002), and $\left\{S^{t}_{0},S^{\theta}_{0}\right\}$ are resolved by Eq. (8). Apart from $s$, frequently mentioned spin parameters are its projection onto the orbital angular momentum $s_{||}$ and another component $s_{\bot}$, which satisfy $s^{2}=s^{2}_{||}+s^{2}_{\bot}$. For resolving the EMRI formation, which will be discussed in the next section, the sBH should have no preference for the direction of the spin, so we focus on the total magnitude $s$ instead of its projections.
@@
LISA and Taji will detect the GW signal with an interferometric measurement of differential optical pathlength modulation along the arms between three free-falling test masses inside spacecraft, i.e. the relative frequency shift $\left\{y_{12}(t),y_{21}(t),y_{13}(t),y_{31}(t),y_{23}(t),y_{32}(t)\right\}$ where the lower indices mark the spacecraft. A change of light-travel distance caused by the GW should be computed by the projection of the above SSB-frame waveforms onto the detector arms, and the response function depends on not only the frequency of signals but also the time due to the typically long duration of EMRI waveforms. Moreover, the laser noise will drown out the GW signal with realistically unequal arms, thus time-delay interferometry (TDI) technology was proposed by Ref. Tinto and Armstrong (1999); Tinto et al. (2002, 2004); Tinto and Dhurandhar (2014). In short, the key point of TDI is the linear combination of time-delay GW projections in order to construct equivalent equal-length arms and suppress the laser noise. For the second-generation TDI designed for evolving length of arms, the Michelson combination, $X$, is given by Tinto et al. (2004)
@@
Other intrinsic parameters are set as $m=10M_{\odot}$, $M=1\times 10^{6}M_{\odot}$, $a=0.99M$, $S^{r}_{0}=S^{\phi}_{0}=0$, $p_{0}=10M$, $e_{0}=0.2$, $\iota_{0}=0.5$, $r_{0}=p_{0}$, $\theta_{0}=\frac{\pi}{2}$ (where the subscript “0” means the initial time), and extrinsic parameters are set as $T=1\text{yr}$, $\theta_{A}=\frac{\pi}{2}$, $\phi_{A}=\frac{\pi}{5}$, $\theta_{N}=\frac{\pi}{3}$, $\phi_{N}=\frac{\pi}{4}$, $\Phi=\frac{\pi}{2}$, $D=1$ Gpc. Note that, although we take initial radial and azimuthal components of $S^{\mu}$ to be 0, they will become nonzero as a result of precession for generic misaligned orbits.
Figure 1: The characteristic strain $h_{c}$ for $s=0.1$ (blue curve) and $s=0.9$ (purple curve), and the noise amplitude $h_{n}$ of Taiji (light gray curve) and LISA (black curve).